# algorithmic

### v2

@@ 3.1 Survival data and notation
Survival analysis, or time-to-event prediction, models the time until an event of interest takes place (Gareth et al., 2021, Ch. 11). We can represent such problems using a two-state model, as depicted in Figure 3(a). In this simplest case, there is one transient state (0, e.g.“alive”) and one absorbing state (1, e.g.“dead”(Andersen and Keiding, 2002). Once an individual enters an absorbing state, no further transitions are possible. The observation for a given instance will in the most simple form consist of a random variable, $T$, representing the time from a given origin ($t=0$) to the occurrence of the event, e.g.“death”. The distribution of $T$ can be characterized by the probability distribution function $F(t)=\text{Pr}(T\leq t)$ or, equivalently, by the survival function $S(t)=1-F(t)=\text{Pr}(T>t)$. Note that $S(t)$ and $F(t)$, respectively, correspond to the probabilities of being in state 0 or 1 at time $t$. In our survival analysis framework, we want to estimate the probability that some specific event occurs at time $T$ later than $t$ (for each $t>0$), i.e., the survival probability, $S\left(t\right)=\text{Pr}(t>T)=1-F\left(t\right)$. In continuous time, the distribution of $T$ can be characterized by the hazard rate function:
@@
In this context, we may define our survival dataset as a set of observations, ($\bm{x}^{(i)}$, $t^{(i)}$, $\delta^{(i)}$), where $\bm{x}^{(i)}\in\mathbb{R}^{d}$ is a feature vector for instance $i$, $t^{(i)}\in\mathbb{R}$ is the time of either censoring or event, depending on which occurred first, and $\delta^{(i)}\in\{0,1\}$ is the “censoring bit”: If $\delta^{(i)}=0$, then $t^{(i)}=c^{(i)}$, where $c^{(i)}$ is the time of censoring, otherwise, if $\delta^{(i)}=1$, then $t^{(i)}=e^{(i)}$, where $e^{(i)}$ is the time of the event. Given survival data $\mathcal{D}=\{(X^{(i)},T^{(i)},\delta^{(i)})\}_{i=1}^{N}$, let $f_{T_{E}\mid X}(\cdot)$ and $F_{T_{E}\mid X}(\cdot)$ represent the conditional density and cumulative distribution functions of the event, respectively, over the event horizon $[0,...,t_{\text{max}}]$. The likelihood of the $i$-th instance experiencing the event of interest can then be determined:
(a) The two-state model with a single transition.
(b) A competing risks model for mortality from two causes.
(c) A multi-event model describing the joint behavior of A and B.
Figure 3: Event models for different scenarios.
@@ 3.2 Multi-state analysis
We now introduce a framework for examining multi-state models called marked point processes (Andersen and Keiding, 2002). A marked point process is a stochastic process $G(t)$, $t\in\mathcal{T}$ with a finite state space $\mathcal{P}=\{0,...,P-1\}$ and with right-continuous sample paths: $G(t+)=G(t)$. Here, $\mathcal{T}=[0,\tau]$ with $\tau\leq\infty$. The process has an initial distribution $\pi_{p}(0)=[\text{Pr}(G(0)=p)$$p\in\mathcal{P}]$. A multi-state process $G(\cdot)$ generates a history $\mathcal{G}_{t}$, which consists of observations of the process in the interval $[0,\tau]$ e.g., Mr. Smith is healthy at the start of the study ($\mathcal{G}_{t=0}$), but becomes ill 10 days later ($\mathcal{G}_{t=10}$), before the study terminates after 100 days ($\tau=100$). Given this history, we can define the transition probabilities by:
$\text{Pr}_{pj}(s,t)=\text{Pr}(G(t)=j\mid G(s)=p,\mathcal{G}_{s^{-}}),$
@@
$\alpha_{pj}(t)=\lim_{\Delta t\to 0}\frac{\text{Pr}_{pj}(t,t+\Delta t)}{\Delta t}.$
@@
$\pi_{p}(t)=\sum_{j\in\mathcal{P}}\pi^{j}(0)\cdot\text{Pr}_{jp}(0,t).$
@@
Competing risks. In competing risks, multiple events (i.e., absorbing states) exist, but only one event (i.e., state transition) can occur for each instance. A competing-risks model thus has one transient state, aka. state 0, and a number of absorbing states, $K$, i.e., event $k$ for $k\in\{1,\ldots,K\}$. Thus, we have $P$ total states and $K=P-1$ event states (see Figure 3(b)). In this model, if an instance is censored before experiencing any events, it simply remained in the transient state and is thus event-free. The transition intensities $\alpha_{0k}(t)$ for $k\in\{1,\ldots,K\}$ are given by the cause-specific hazard functions:
$\alpha_{k}(t)=\lim_{\Delta t\to 0}\frac{{}\text{Pr}(\text{Dead from cause}\;k\;\text{at}\;t+\Delta t\mid T\geq t)}{\Delta t}\text{,}$
@@
$\text{Pr}_{00}(0,t)=S(t)=\text{Pr}(T>t)=\exp\left(-\int_{0}^{t}\sum_{k=1}^{K}\alpha_{k}(u)\,du\right)\text{,}$
@@
Multi-event. In multi-event survival analysis, we can observe multiple events for each instance $i$, thus they are not mutually-exclusive. Figure 3(c) exhibits a multi-event model, which describes the behavior of two events $A$ and $B$, and their joint event, $AB$. As an example, consider the probability that Mr. Smith will have the onset of pneumonia (A) and COVID-19 (B) at the same time. If the time until the onset of $B$ does not depend on whether the patient has $A$ – that is, $\alpha_{0B}=\alpha_{A,AB}$ – then $A$ is called locally dependent on $B$. If the onset of $A$ is different depending on whether the patient has $B$ – that is, $\alpha_{0A}\neq\alpha_{B,AB}$ – then $B$ is not locally dependent on $A$. Here, this temporal sequence of events thus provides asymmetric dependence between events (Andersen and Keiding, 2002).
@@ 4.1 Preliminaries
We now extend the previous single-event notation to a multi-event notation over $P$ states. Let $\mathcal{D}=\left\{\left(\bm{x}^{(i)},\bm{t}^{(i)},\bm{\delta}^{(i)}\right)\right\}_{i=1}^{N}$ be the dataset, where $\bm{x}^{(i)}\in\mathbb{R}^{d}$ is a $d$-dimensional vector of features, $\bm{t}^{(i)}\in\{1,2,\ldots,T_{\text{max}}\}^{K}$ is a vector of observed times for the $i$-th person now across $K$ multiple events, each at some time, and $\bm{\delta}^{(i)}\in\{0,1\}^{K}$ is vector of event indicators for each event. Moreover, let $e_{k}^{(i)}\in\{1,2,\ldots,T_{\text{max}}\}$ denote the event time and let $c_{k}^{(i)}\in\{1,2,\ldots,T_{\text{max}}\}$ denote the censoring time for the $i$-th instance for event $k$ over the event horizon, thus $t_{k}^{(i)}=e_{k}^{(i)}$ if $\delta_{k}^{(i)}=1$, otherwise, $t_{k}^{(i)}=c_{k}^{(i)}$ if $\delta_{k}^{(i)}=0$. We define a multi-event survival problem as predicting the probability of instance $i$ experiencing event $k$ at each time point $t\in\{1,2,...,T_{\text{max}}\}$. This produces a matrix of survival functions $\bm{S}^{(i)}=[\bm{s}_{1}^{(i)},\bm{s}_{2}^{(i)},\ldots,\bm{s}_{P}^{(i)}]$ for instance $i$, where $P$ is the number of states. Recall that the survival function is the probability that some event occurs at time $T$ later than $t$i.e.$\bm{s}_{p}^{(i)}=[\text{Pr}(e_{p}^{(i)}>1),\text{Pr}(e_{p}^{(i)}>2),\ldots,\text{Pr}(e_{p}^{(i)}>T_{\text{max}})]$ describes the probability that instance $i$ does not transition to state $p$ until $t$ for $t\in\{1,2,...,T_{\text{max}}\}$.
@@
Let us revisit our fictive patient, Mr. Smith. participates in the aforementioned clinical trial and enrolls in the study at baseline ($t=0$) with his covariates $\bm{x}^{(i)}$. After taking the drug, Mr. Smith experienced nausea after 10 days and fatigue after 20 days, and was never censored for any events (i.e., we observed all events). Using the proposed multi-state notation, we would encode Mr. Smith’s event information as follows:
@@ 4.2 Mixture model architecture
$f(t;\,\beta,\eta)$ $=\ \frac{\eta}{\beta}\left(\frac{t}{\beta}\right)^{\eta-1}\exp\left({-\left(\frac{t}{\beta}\right)^{\eta}}\right),$
$F(t;\,\beta,\eta)$ = $=\ 1-\exp\left({-\left(\frac{t}{\beta}\right)^{\eta}}\right),$
@@
Here, $x^{(i)}\in\mathbb{R}$ is the input vector of feature values describing the $i$-th instance, and $\Phi_{\theta}(\cdot)$ is a feature representation layer, which is shared among events. This layer accounts for inter-event dependencies captured in the features. The other terms are specific to each output event. The parameters $\tilde{\beta}_{\psi}\in\mathbb{R}$ and $\tilde{\eta}_{\psi}\in\mathbb{R}$ are learnable log-scale biases that initialize the log-transformed scale and shape parameters, respectively, for the $\psi$-th component. $\text{act}(\cdot)$ is a non-linear activation function, in our case SELU (Klambauer et al., 2017). Finally, $\zeta_{\psi}$ and $\xi_{\psi}$ are learnable weight vectors for the scale and shape parameters, respectively, of the $\psi$ Weibull distribution. As a result, we can define the PDF and CDF for each marginal as:
@@ 4.3 Competing-risks likelihood
Given survival data $\mathcal{D}=\{(X^{(i)},T^{(i)},\delta^{(i)})\}_{i=1}^{N}$, recall that $f_{T_{E}\mid X}(\cdot)$ represents the conditional density function, where $T_{E}$ is the event distribution. In order to estimate Eqs. (2) and (3) using MLE, we adopt the likelihood in Eq. (1) to the competing-risks scenario under the conditional independence assumption, now using the survival function $S_{T\mid X}(\cdot)$, which represents the likelihood that the event (or censoring) will take place after the argument time $t$. This approach mimics the one in DSM (Nagpal et al., 2021), however, we also consider the event-free state (i.e., the transient state) as an event in the likelihood function by calculating the likelihood across $P$ states instead of $K$ events. Figure 6 shows an example of the likelihood composition based on the previous competing-risks model in Figure 3(b).
@@
This explicitly models the probability of staying event-free (i.e., not experiencing any of $K$ events) for instance $i$ over the event horizon $[0,...,t_{\text{max}}]$, and thus allows us to “predict” the event-free (transient) distribution, i.e., will the instance $i$ remain event-free over the follow-up period or experience one of $K$ events. Our approach also differs from Nagpal et al. (2021) as we optimize the log-likelihood directly instead of deriving a lower bound on it. Given the survival dataset $\mathcal{D}$ with $N$ instances and $K$ competing risks, where $\delta_{j}^{(i)}$ indicates the event indicator for $j$th state for the $i$th instance, we can write the composite likelihood as follows (see also Figure 6):
$\begin{split}\mathcal{L}_{\text{CR}}(\mathcal{D})\ =\ \prod_{i=1}^{N}\prod_{j=1}^{P}&\,f_{T_{E_{j}}\mid X}(T_{\text{obs}}^{(i)}\mid X^{(i)})^{\delta^{(i)}_{j}}\cdot S_{T_{E_{j}}\mid X}(T_{\text{obs}}^{(i)}\mid X^{(i)})^{1-\delta^{(i)}_{j}}.\end{split}$
@@ 4.4 Multi-event likelihood with trajectories
We now extend our model to the multi-event scenario, where events (i.e., state transitions) can happen in any order and at any time, and are not necessarily mutually-exclusive. To this end, we can simply extend Eq. 6 to a multi-state scenario as a composite likelihood function over the $P$ states, with different observed times for the different state transitions, omitting the log-operation for brevity:
@@
which differs from Eq. 6 only by using $T^{(i)}_{\text{obs}_{(j)}}$, which now indicates the observed time for $j$th state transition – i.e., instance $i$ was last observed in state $j$ at time $T^{(i)}_{\text{obs}_{(j)}}$. This likelihood mimics the likelihood in Tjandra et al. (2021), who calculated the event-specific likelihood over $K$ events, not considering the event-free state nor the temporal ordering between events. However, many real-world events do not happen in isolation, but have some relationship to each other in time.
An example is the illness-death model (Fix and Neyman, 1951), where patients start out healthy (state 0) and then may become ill (event A) and go on to die (event B). In general, there are four possible scenarios for two events at any specific time: (1) Neither event A nor B has happened (i.e., both are censored), (2) only event A happened (but B is censored), (3) only event B has happened (but A is censored), and (4) both events happened, and event A occurred before event B, i.e., the patient became ill and then died. In the last case, we have a known ordering of events, assuming that patients who experience both A and B must first become ill and then die. (Note that we do not consider the possibility of people dying before the onset of a disease in this example).
@@
$\mathcal{L}_{\text{total}}(\mathcal{D})$ $=\frac{1}{N}\mathcal{L}_{\text{ME}}(\mathcal{D})+\frac{1}{N}\mathcal{L}_{\text{trajectory}}(\mathcal{D}).$

### v3

@@ 3.1 Survival data and notation
Survival analysis, or time-to-event prediction, models the time until an event of interest takes place [13, Ch. 11]. We can represent such problems using a two-state model (see Figure 3). In this simplest case, there is one transient state (0, e.g.“alive”) and one absorbing state (1, e.g.“dead”[2]. Once an individual enters an absorbing state, no further transitions are possible. The observation for a given instance will in the most simple form consist of a random variable, $T$, representing the time from a given origin ($t=0$) to the occurrence of the event, e.g.“death”. The distribution of $T$ can be characterized by the probability distribution function $F(t)=\text{Pr}(T\leq t)$ or, equivalently, by the survival function $S(t)=1-F(t)=\text{Pr}(T>t)$. Note that $S(t)$ and $F(t)$, respectively, correspond to the probabilities of being in state 0 or 1 at time $t$. In our survival analysis framework, we want to estimate the probability that some specific event occurs at time $T$ later than $t$ (for each $t>0$), i.e., the survival probability, $S\left(t\right)=\text{Pr}(t>T)=1-F\left(t\right)$. In continuous time, the distribution of $T$ can be characterized by the hazard rate function:
$a\left(t\right)=-d\log S(t)/{dt}=\lim_{\Delta t\rightarrow 0}\frac{\text{Pr}\left(t<T\leq t+\Delta t|T>t\right)}{\Delta t}\text{,}$ (1)
where $S(t)$ is the survival function:
@@
Figure 3: A single-event model with a single transition.
Figure 4: A competing-risks model for mortality from two causes.
Figure 5: A multi-event model describing the joint behavior of A and B.
@@ 3.2 Multi-state analysis
We now introduce a framework for examining multi-state models called marked point processes [2]. A marked point process is a stochastic process $G(t)$, $t\in\mathcal{T}$ with a finite state space $\mathcal{P}=\{0,...,P-1\}$ and with right-continuous sample paths: $G(t+)=G(t)$. Here, $\mathcal{T}=[0,\tau]$ with $\tau\leq\infty$. The process has an initial distribution $\pi_{p}(0)=[\text{Pr}(G(0)=p)$$p\in\mathcal{P}]$. A multi-state process $G(\cdot)$ generates a history $\mathcal{G}_{t}$, which consists of observations of the process in the interval $[0,\tau]$e.g., Mr. Smith is healthy at the start of the study ($\mathcal{G}_{t=0}$), but becomes ill 10 days in ($\mathcal{G}_{t=10}$), before the study terminates after 100 days ($\tau=100$). Given this history, we can define the transition probabilities by:
$\text{Pr}_{pj}(s,t)=\text{Pr}(G(t)=j\mid G(s)=p,\mathcal{G}_{s^{-}}),$ (3)
@@
$\alpha_{pj}(t)=\lim_{\Delta t\to 0}\frac{\text{Pr}_{pj}(t,t+\Delta t)}{\Delta t}.$ (4)
@@
$\pi_{p}(t)=\sum_{j\in\mathcal{P}}\pi^{j}(0)\cdot\text{Pr}_{jp}(0,t).$ (5)
@@
Competing risks. In competing risks, multiple events (i.e., absorbing states) exist, but only one event (i.e., state transition) can occur for each instance. A competing-risks model thus has one transient state, aka. state 0, and a number of absorbing states, $K$, i.e., event $k$ for $k\in\{1,\ldots,K\}$. Thus, we have $P$ total states and $K=P-1$ event states (see Figure 4). In this model, if an instance is censored before experiencing any events, it simply remained in the transient state and is thus event-free. The transition intensities $\alpha_{0k}(t)$ for $k\in\{1,\ldots,K\}$ are given by the cause-specific hazard functions:
$\alpha_{k}(t)=\lim_{\Delta t\to 0}\frac{{}\text{Pr}(\text{Dead from cause}\;k\;\text{at}\;t+\Delta t\mid T\geq t)}{\Delta t}\text{,}$ (6)
@@
$\text{Pr}_{00}(0,t)=S(t)=\text{Pr}(T>t)=\exp\left(-\int_{0}^{t}\sum_{k=1}^{K}\alpha_{k}(u)\,du\right)\text{,}$ (7)
@@
Multi-event. In multi-event survival analysis, we can observe multiple events for each instance $i$, thus they are not mutually exclusive (see Figure 5). Such a model can, for example, describe the behavior of two events $A$ and $B$, and their joint event, $AB$. Consider the probability that a patient will have the onset of pneumonia (A) and COVID-19 (B) at the same time. If the time until the onset of $B$ does not depend on whether the patient has $A$ – that is, $\alpha_{0B}=\alpha_{A,AB}$ – then $A$ is called locally dependent on $B$. If the onset of $A$ is different depending on whether the patient has $B$ – that is, $\alpha_{0A}\neq\alpha_{B,AB}$ – then $B$ is not locally dependent on $A$.
@@ 4.1 Preliminaries
We now extend the previous single-event notation to a multi-event notation over $P$ states. Let $\mathcal{D}=\left\{\left(\bm{x}^{(i)},\bm{t}^{(i)},\bm{\delta}^{(i)}\right)\right\}_{i=1}^{N}$ be the dataset, where $\bm{x}^{(i)}\in\mathbb{R}^{d}$ is a $d$-dimensional vector of features, $\bm{t}^{(i)}\in\{1,2,\ldots,T_{\text{max}}\}^{K}$ is a vector of observed times for the $i$-th instance now across $K$ multiple events, each at some time, and $\bm{\delta}^{(i)}\in\{0,1\}^{K}$ is vector of event indicators for each event. Moreover, let $e_{k}^{(i)}\in\{1,2,\ldots,T_{\text{max}}\}$ denote the event time and let $c_{k}^{(i)}\in\{1,2,\ldots,T_{\text{max}}\}$ denote the censoring time for the $i$-th instance for event $k$ over the event horizon, thus $t_{k}^{(i)}=e_{k}^{(i)}$ if $\delta_{k}^{(i)}=1$, otherwise, $t_{k}^{(i)}=c_{k}^{(i)}$ if $\delta_{k}^{(i)}=0$. We define a multi-event survival problem as predicting the probability of instance $i$ experiencing event $k$ at each time point $t\in\{1,2,...,T_{\text{max}}\}$. This produces a matrix of survival functions $\bm{S}^{(i)}=[\bm{s}_{1}^{(i)},\bm{s}_{2}^{(i)},\ldots,\bm{s}_{P}^{(i)}]$ for instance $i$, where $P$ is the number of states. Recall that the survival function is the probability that some event occurs at time $T$ later than $t$i.e.$\bm{s}_{p}^{(i)}=[\text{Pr}(e_{p}^{(i)}>1),\text{Pr}(e_{p}^{(i)}>2),\ldots,\text{Pr}(e_{p}^{(i)}>T_{\text{max}})]$ describes the probability that instance $i$ does not transition to state $p$ until $t$ for $t\in\{1,2,...,T_{\text{max}}\}$.
@@
Let us revisit our fictive patient, Mr. Smith. He participates in the aforementioned clinical trial and enrolls in the study at baseline ($t=0$) with his covariates $\bm{x}^{(i)}$. After taking the drug, Mr. Smith experienced nausea after 10 days and fatigue after 20 days, and was never censored for any events (i.e., we observed all events). Using the proposed multi-state notation, we would encode Mr. Smith’s event information as follows:
@@ 4.2 Mixture model architecture
$f(t;\,\beta,\eta)$ $=\ \frac{\eta}{\beta}\left(\frac{t}{\beta}\right)^{\eta-1}\exp\left({-\left(\frac{t}{\beta}\right)^{\eta}}\right),$ (9)
$F(t;\,\beta,\eta)$ = $=\ 1-\exp\left({-\left(\frac{t}{\beta}\right)^{\eta}}\right),$ (10)
@@
Here, $x^{(i)}\in\mathbb{R}$ is the input vector of feature values describing the $i$-th instance, and $\Phi_{\theta}(\cdot)$ is a feature representation layer, which is shared among events. This layer accounts for inter-event dependencies captured in the features. The other terms are specific to each output event. The parameters $\tilde{\beta}_{\psi}\in\mathbb{R}$ and $\tilde{\eta}_{\psi}\in\mathbb{R}$ are learnable log-scale biases that initialize the log-transformed scale and shape parameters, respectively, for the $\psi$-th component. $\text{act}(\cdot)$ is a non-linear activation function, in our case SELU [26]. Finally, $\zeta_{\psi}$ and $\xi_{\psi}$ are learnable weight vectors for the scale and shape parameters, respectively, of the $\psi$ Weibull distribution. As a result, we can define the PDF and CDF for each marginal as:
@@ 4.3 Competing-risks likelihood
Given survival data $\mathcal{D}=\{(X^{(i)},T^{(i)},\delta^{(i)})\}_{i=1}^{N}$, let $f_{T\mid X}(\cdot)$ and $F_{T\mid X}(\cdot)$ represent the conditional density and cumulative distribution functions of an outcome of interest (e.g., event or censoring) over the event horizon $[0,...,t_{\text{max}}]$. The likelihood of the $i$-th instance experiencing the event of interest can then be determined:
In order to estimate Eqs. (11) and (12) using MLE, we adopt the likelihood in Eq. (15) to the competing-risks scenario under the conditional independence assumption, now using the survival function $S_{T\mid X}(\cdot)$, which represents the likelihood that the event (or censoring) will take place after the argument time $t$. This approach mimics the one in DSM [36], however, we also consider the event-free state (i.e., the transient state) as an event in the likelihood function by calculating the likelihood across $P$ states instead of $K$ events. Figure 8 shows an example of the likelihood composition based on the previous competing-risks model (see Figure 4).
@@
This explicitly models the probability of staying event-free (i.e., not experiencing any of $K$ events) for instance $i$ over the event horizon $[0,...,t_{\text{max}}]$, and thus allows us to “predict” the event-free (transient) distribution, i.e., will the instance $i$ remain event-free over the follow-up period or experience one of $K$ events. Our approach also differs from Nagpal et al. [36] as we optimize the log-likelihood directly instead of deriving a lower bound on it. Given a survival dataset $\mathcal{D}$ with $N$ instances and $K$ competing risks, where $\delta_{j}^{(i)}$ denotes the event indicator for $j$th state for the $i$th instance, we can write the composite likelihood as follows (see also Figure 8):
$\begin{split}\mathcal{L}_{\text{CR}}(\mathcal{D})\ =\ \prod_{i=1}^{N}\prod_{j=1}^{P}&\,f_{T_{E_{j}}\mid X}(T_{\text{obs}}^{(i)}\mid X^{(i)})^{\delta^{(i)}_{j}}\cdot S_{T_{E_{j}}\mid X}(T_{\text{obs}}^{(i)}\mid X^{(i)})^{1-\delta^{(i)}_{j}}.\end{split}$ (16)
@@ 4.4 Multi-event likelihood with trajectories
We now extend our model to the multi-event scenario, where events (i.e., state transitions) can happen in any order and at any time, and are not necessarily mutually exclusive. To this end, we can simply extend Eq. 17 to a multi-state scenario as a composite likelihood function over the $P$ states, with different observed times for the different state transitions, omitting the log-operation for brevity:
@@
which differs from Eq. 17 only by using $T^{(i)}_{\text{obs}_{(j)}}$, which now indicates the observed time for $j$th state transition – i.e., instance $i$ was last observed in state $j$ at time $T^{(i)}_{\text{obs}_{(j)}}$. This likelihood mimics the likelihood in Tjandra et al. [51], who calculated the event-specific likelihood over $K$ events, not considering the event-free state nor the temporal ordering between events. However, many real-world events do not happen in isolation, but have some relationship to each other in time.
An example is the illness-death model [11], where patients start out healthy (state 0) and then may become ill (event A) and go on to die (event B). In general, there are four possible scenarios for two events at any specific time (see Table 1): (1) Neither event A nor B has happened (i.e., both are censored), (2) only event A happened (but B is censored), (3) only event B has happened (but A is censored), and (4) both events happened, and event A occurred before event B, i.e., the patient became ill and then died. In the last case, we have a known ordering of events, assuming that patients who experience both A and B must first become ill and then die. (Note that we do not consider the possibility of people dying before the onset of a disease in this example).
@@
$\mathcal{L}_{\text{total}}(\mathcal{D})$ $=\frac{1}{N}\mathcal{L}_{\text{ME}}(\mathcal{D})+\frac{1}{N}\mathcal{L}_{\text{trajectory}}(\mathcal{D}).$ (20)